A roulette wheel has 38 outcomes labeled 1 through 36 plus 0 and 00. The wheels are supposed to be designed so that each outcome is equally likely. The numbers 0 and 00 are often referred to as house numbers because the only way that a player can win when these outcomes are observed is by directly betting on the numbers. A great deal of money wagered on a roulette wheel is wagered on odd or even numbers, or columns or rows of numbers. The numbers 0 and 00 are not in any row or column, nor are they odd or even. Write your answers as fractions!

a. What is the probability of observing an even number?

  1. What is the probability of observing a number between 1 and 12, inclusive?
Answer #1

a)

probability of observing an even number =18/38 =9/19 (since there are 18 even number out of 38)

b)

probability of observing a number between 1 and 12, inclusive =12/38 =6/19 (sicne these are 12 number out of total 38 possible outcomes)